OpenStudy (anonymous):

find the root of this equation: 0=2x+8

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

like move the 8?

OpenStudy (anonymous):

Like we need to find out what x equal yes

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

So show me what you have?

OpenStudy (anonymous):

okay thx=3

OpenStudy (anonymous):

okay thx=3

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

nope \[0=2x+8\] Subtract 8 from both sides:\[-8=2x+8-8 \to -8=2x\] Divide both sides by 2: \[\frac{-8}{2} =\frac{2x}{2} \to -4=x\]

OpenStudy (anonymous):

Yup you got it

OpenStudy (anonymous):

thx